### v1

@@ 2.3 Testing General Relativity
X-ray reflection spectroscopy is currently the technique that can provide the most stringent test of the Kerr hypothesis. We simulated a 30 ks observation of a bright black hole X-ray binary with eXTP/SFA and NICER/XTI assuming that the compact object is a Kerr black hole with spin parameter $a_{*}=0.5$ and that the inclination angle of the disk is $i=70$ deg. We fit the data with the Johannsen metric allowing for a non-vanishing deformation parameter $\alpha_{13}$. For the simulations and the fits, we used the model relxill_nk [121, 122], which is an extension of the relxill package to non-Kerr spacetimes. The constraints on the black hole spin parameter $a_{*}$ and the deformation parameter $\alpha_{13}$ (where $\alpha_{13}=0$ corresponds to the Kerr metric) are shown in Fig. 11 (after marginalizing over all other parameters): the eXTP/SFA measurements are in red and the NICER/XTI measurements are in blue. eXTP can roughly halve the uncertainty of the NICER measurement with the same exposure time. Fig. 11 also shows the eXTP/SFA measurements of a putative 300 ks observation of the same source (in green). A longer exposure time can clearly improve the constraints on possible deviations from the Kerr geometry. Unlike spin measurements, where we want to measure the spin of as many black holes as possible, in the case of tests of General Relativity we can properly select the most suitable sources, where systematic uncertainties are under control and subdominant with respect to the statistical ones. It is indeed plausible to image that either General Relativity works or does not, regardless of the source.
@@ 3 Exploring Accretion Physics in Extreme Gravity
A great number of efforts have been made towards establishing the intrinsic connection between black hole X-ray binaries and AGN, such as the fundamental plane of black hole activity and the unified scheme for black hole accretion, in which it is proposed that different types of AGN are determined by Eddington ratio (Eddington rate scaled mass accretion rate) and the viewing angle. As that of in black hole X-ray binaries, although the basic picture has been there, many key questions are still needed to be settled, such as, the structure of the disk and corona, the physical mechanism of the production and evolution of QPOs, the origin of quasi-periodic eruptions (QPEs), AGN feedback etc.
The process of accretion onto black holes plays an important role in shaping the Universe. The energy released in a small region (e.g., a few tens of gravitational radii) near the black holes in AGN can cause significant feedback which affects the evolution of their host galaxies.
3.1 The geometry and dynamics of the accretion flow in stellar mass BHs
@@ 3.1.1 Disk-corona time delays
In addition to spectral analysis, timing analysis has been employed to study the accretion geometry and its evolution. Low- and high-frequency noise in the power spectrum of the X-ray fast variations carries information about the accretion flow [133]. In some X-ray binaries, it has been observed that at low frequencies, the hard X-ray variation lags behind the soft X-ray (hard delay), and the time delay decreases with decreasing optical variation time scale. These ‘hard lags’ are thought to arise from perturbations in the accretion rate at different radii of the accretion disk propagating inward [134]. In contrast, at high variability frequencies (typically $>1$ Hz), soft X-ray variations are observed to lag behind those in hard X-rays (soft delay or ‘soft lags’). This effect may originate from the time difference corresponding to the irradiation of the accretion disk by the thermal accretion flow, i.e., the X-ray reverberation signal.
@@
Figure 12: Cross-section of the hyperbolic corona in the black hole frame. a and b are the semi-major and semi-minor axis of the hyperboloid. For disk-like corona (black solid line): a = 10, b = 2, upper boundary height $\textit{h}_{\rm t}=30$, lower boundary height $\textit{h}_{\rm b}=6$. For jet-like corona (red solid line): a = 120, b = 200, $\textit{h}_{\rm t}=180$, $\textit{h}_{\rm b}=2$. The units of all numbers are $R_{\rm g}$.
@@
$\left\{\begin{array}[]{cl}\frac{x^{2}+y^{2}}{a^{2}}-\frac{z^{2}}{b^{2}}=1\\ h_{\rm b}\leq|z|\leq h_{\rm t}\\ \end{array}\right.$ (2)
where a and b are the semi-major and semi-minor axis of the hyperboloid. $\textit{h}_{\rm t}$ and $\textit{h}_{\rm b}$ represent the heights of the top and bottom boundaries, respectively. The reference coordinate system is Cartesian, with the black hole at zero, where a geometrically thin and optically thick disk is situated in the xy plane, depicted by the blue solid line in Fig. 12. The disk has a truncation radius of $10R_{\rm g}$ and an external radius of $500R_{\rm g}$. Two distinct geometric types (disk-like and jet-like) correspond to varying parameters of the hyperboloid (see Fig. 12). The disk-like corona is wider than it is tall, resembling a radially extended disk, while the jet-like corona appears the opposite, resembling a jet.
@@
For the light curve, the count rates in each energy bin were multiplied by the mean effective area for the corresponding energy bin. The count rates within the specified energy bands (0.5-1.0 keV and 2.0-5.0 keV) were summed to produce the light curve. Using the Python library Stingray v2.0.0 [142, 143, 144], the time lag-frequency between the 0.5-1.0 keV and 2.0-5.0 keV bands was computed. Fig.  14 presents the results, where a positive lag indicates a hard lag, meaning the hard photons are delayed relative to the soft photons. The simulation results in Fig. 13 and Fig. 14 clearly indicate that, while the energy spectra of the disk-like and jet-like corona are indistinguishable, the frequency-dependent lag reveals significant differences.
@@ 3.1.2 Reflection reverberation
Figure 15: Spectral model and simulated eXTP-SFA lag vs. energy spectra for a 1 Crab hard state black hole XRB, showing that two distinct inner geometries for disk inner radius ($r_{\rm in}$ and source height ($h$) can be easily distinguished with the lags. The upper panel shows the assumed spectral models with disk emission (dotted lines) set to have similar disk temperature, making the overall spectral energy distributions nearly identical. The lower panel shows the simulated lag data. The simulation assumes 100 ks exposures and shows the lags obtained for the 10-20 Hz frequency range, where evidence for soft reverberation lags has been seen by NICER and where reverberation delays are expected to dominate over propagation delays, which are washed out on these time-scales. The simulation uses the reltrans [71, 72] lamp-post model for reverberation lags. The lag model curves are shown with the smooth solid lines and data and model are set to zero lag at 2 keV so the shapes can be easily compared.
@@
Figure 16: Simulated eXTP/PFA data for a generic BH X-ray binary source at the initial stages of an outburst. The binary is assumed to have an inclination of $i=30^{\circ}$, with an incident spectrum characterized by a photon index $\Gamma=1.6$ in the 2–8 keV range and an unabsorbed flux of 100 mCrab. The model PD corresponds to a slab corona geometry with an outflow relative velocity of $\beta=0$ (static, black solid line) and $\beta=0.4$ (mildly relativistic, black dashed line), based on [88]. An eXTP exposure of 100 ksec is sufficient to distinguish between these alternatives.
@@ 3.1.3 Polarization measurements of coronal geometry
These questions are not expected to be solved with the IXPE instrument, given the fast evolution of the transient black hole X-ray binaries, their typically dim fluxes in the initial outburst stages and the requirement of polarization measurement with sub-per cent accuracy. To get to a precision of 0.2% on polarization degree, required to constrain the outflow parameters, a generic 100 mCrab hard state black hole X-ray binary source would need to be observed by IXPE for more than a week (accounting for the $\sim$50effective exposure time). This time is comparable to the characteristic timescales of the outburst evolution, preventing precise measurements of the geometry at distinct times, and a clear understanding of geometric evolution. In contrast, as shown in Fig. 16, eXTP/PFA brings the exposure time down to about a day, hence becoming uniquely suited to track the evolution of accretion geometry for black hole transients.
@@ 3.1.4 Low-frequency QPOs as a disk-corona probe
The geometry of the corona surrounding black holes is not static during outbursts. Recent studies propose that the coronal geometry undergoes dynamic evolution, particularly during spectral state transitions from the hard to soft state [e.g. 150, 151]. High-cadence observations with the eXTP mission will be uniquely positioned to probe this evolution through its unprecedented combination of simultaneous X-ray spectral, timing, and polarimetric measurements. Furthermore, eXTP’s capacity for prolonged exposure significantly enhances the likelihood of capturing rapid changes of corona during state transitions [e.g. 152, 153, 154].
@@
3.2 Mapping the inner regions of SMBHs
@@ 3.2.1 X-ray reverberation mapping of the inner regions
As for stellar-mass black hole accreting systems (XRBs), the measurement of energy-dependent time delays, i.e., lag-energy spectra, are useful to test detailed models for the innermost emitting regions of AGN, such as propagating fluctuations and reverberation [e.g., 146]. However, the spectral-timing measurements require observations with very high count rates, so can only be done for a handful of bright AGNs with current instruments. Furthermore, for past and current proportional-counter instruments with large collecting area at Fe K energies (e.g. RXTE, ASTROSAT), the spectral resolution is poor, which prevents detailed investigation of reverberation properties in the Fe K band. With its large effective area and CCD-quality spectral-resolution in the 0.5–10 keV band, eXTP/SFA will provide much better signal-to-noise measurements of the lag-energy spectra, at sufficient resolution to resolve Fe K reverberation signatures.
@@
Given the large effective area of eXTP-SFA, we can perform X-ray reverberation mapping of both the continuum and Fe K line for the brightest nearby intermediate-mass black holes (IMBHs, 102-106 M), such as NGC 4395. IMBHs serve as a crucial population bridging stellar-mass black holes and supermassive black holes (SMBHs). Measuring X-ray time lags in IMBHs is essential for filling the gap between these two mass regimes and for establishing a global scaling relation between the continuum/emission-line lags and black hole mass [146]. These measurements also provide key insights into the corona-disk-jet connection across different accretion modes [179]. Furthermore, recent IXPE observations [180] suggest that the corona geometry extends along the same direction as the accretion disk, challenging the simple lamppost model assumption. eXTP offers a unique opportunity to test this corona illumination and disk reprocessing scenario, providing new constraints on the physics of accretion and black hole environments.
@@ 3.2.2 AGN non-thermal superflare from an orbiting hotspot at the SMBH inner radii
Recently, Very Large Telescope Interferometer (VLTI) observed several near-infrared superflares which might be from hot spots moving around Sgr A* at only 7–10 gravitational radii from the central black hole (Gravity Collaboration et al. 2018; Gutierrez, Nemmen & Cafardo 2020). It is believed that these hot spots are likely due to relativistic magnetic reconnection events and they might be also the sources of X-ray flares which exhibit a double-peak structure in the in the light curve of the total X-ay flux due to their orbital motion. Here we simulate the X-ray light curve of these hot spots observed by eXTP-SFA.
@@
Figure 18 (left) shows a simulation of the light curve produced by a hot spot orbiting a black hole for one period, as observed by eXTP-SFA in the 2–10 keV energy band. We simulate a 4.2 ks observation with an average flux of $10^{-11}$ erg cm-2 s-1. The black hole mass is assumed to $4.14\times 10^{6}\,M_{\odot}$ and black hole spin is 0.998. The orbital radius of the hot spot is $r_{\mathrm{spot}}=10\,R_{g}$. Its radiation is assumed as a power law spectrum with index $\Gamma=2$. The observing inclination is assumed to be $75^{\circ}$. There is an obvious double-peak structure in the light curve, with a time interval of 500 s between the two peaks. In the lower subpanel, we show the ratio between the data and the best fit model. We obtain a good fit and recover the input parameters.
Figure 18 (right) shows the error contours (1, 2, and 3 $\sigma$ in red, green and blue, respectively) of hot spot orbital radius and observing inclination. The orbital radius of the hot spot can be measured to a precision $<3\,R_{g}$ at $3\sigma$ confidence, while the inclination is measured to a precision of $<1^{\circ}$. The technique also provides us with a way to measure the black hole spin, if the mass of the black hole is measured independently (e.g. via optical reverberation mapping [184]). This would apply to any hot spot radius and hence is independent of whether or not the disk is assumed to extend to the ISCO, as is the case for spin constraints from spectral fitting. Thanks to its unprecedented throughput, eXTP will be able to perform such measurements for any AGN with X-ray flux above 1 mCrab.
@@ 3.2.3 Quasi Periodic Oscillations from SMBHs
AGN, powered by accretion on to supermassive black holes (SMBHs) at the center of of galaxies, are thought to be scaled-up versions of black hole XRBs [185, 186, 187, 188]. A compelling line of evidence for this postulation is the striking similarity in the variability of the X-ray radiation between AGN and BHXRBs [e.g. 189, 190, 191, 192, 193]. The power spectral densities (PSDs) of both can be well described as red-noise. However, the characteristic QPOs, which have been observed in the X-ray light curves of dozens of BHXRBs [e.g. 194, 127], are rarely detected in AGN. More detections of QPOs in SMBHs will strengthen the connection between AGN and BHXRBs. Also, the properties of QPOs in AGN can provide a powerful probe to diagnose the inner structure of the X-ray emitting region around SMBHs.
@@
So far, only a few possible QPOs have been found in accreting systems with SMBHs: the Narrow-Line Seyfert 1 (NLS1) galaxies RE J1034+396 [201], 1H 0707-495 [202], ESO 113-G010 [203], Mrk 766 [204], MCG-06-30-15 [205], MS 2254.9–3712 [206], the ultrasoft AGN candidate 2XMM J123103.2+110648, as well as the tidal disruption events (TDE) Swift J164449.3+573451 [207], etc. But the reliability of some cases are still under debate [e.g. 208]. Considering the the dependence of the timescale on the BH mass [127, 209], it is not hard to understand the lack of QPO detections in SMBHs. More QPOs in SMBHs will be found if more high-quality observations with long, continuous, uninterrupted exposure can be taken.
Among all the QPOs detected in SMBHs, the most acceptable and significant case is the one in RE J1034+396. The continuum of the PSD can be fitted by a powerlaw model, and the QPO signal around $2.7\times 10^{-4}$ Hz can be well described by a Lorentzian profile, as shown in the left panel of Figure 19. The 0.3-10 keV count rate (PN+MOS1+MOS2) of RE J1034+396 observed with XMM-Newton is about 7.2 counts s-1, which corresponds to a flux of about $1\times 10^{-11}\,\rm ergs\,s^{-1}\,cm^{-2}$, by assuming a typical power-law spectrum model with photon index of 2 and hydrogen column density of $3\times 10^{20}\,\rm cm^{-2}$. Using the RE J1034+396 PSD model as input, and the count rate of about 20 counts s-1, we simulate 1000 lightcurves of RE J1034+396 observed with SFA, with duration of 140 ks. In over 80% of the observations, significant QPO signals with confidence level higher than 99.99% can be detected. The detection probability will decrease to about 50% when the flux of source is about $2.5\times 10^{-13}\,\rm ergs\,s^{-1}\,cm^{-2}$. The capability of SFA can expand the sample of QPOs in SMBHs. Or the observations can help answer another question: whether the QPOs are really rarer in accreting systems with SMBHs than those with stellar BHs.
Figure 19: Left panel: the PSD of RE J1034+396 observed with XMM-Newton. The blue line represents the best-fit power law, the orange line represents the PSD model with an added Lorentzian profile, and the red dashed line represents the 99.99% confidence level. Middle and right panels: the examples of simulated PSDs observed with SFA, with flux of $1\times 10^{-11}\,\rm ergs\,s^{-1}\,cm^{-2}$ and $2.5\times 10^{-13}\,\rm ergs\,s^{-1}\,cm^{-2}$, respectively.

### v2

@@ 2.3 Testing GR
X-ray reflection spectroscopy is currently the technique that can provide the most stringent test of the Kerr hypothesis. We simulated a 30 ks observation of a bright black hole X-ray binary with eXTP/SFA and NICER/XTI assuming that the compact object is a Kerr black hole with spin parameter $a_{*}=0.5$ and that the inclination angle of the disk is $i=70$ deg. We fit the data with the Johannsen metric allowing for a non-vanishing deformation parameter $\alpha_{13}$. For the simulations and the fits, we used the model relxill_nk [142, 143], which is an extension of the relxill package to non-Kerr spacetimes. The constraints on the black hole spin parameter $a_{*}$ and the deformation parameter $\alpha_{13}$ (where $\alpha_{13}=0$ corresponds to the Kerr metric) are shown in Fig. 11 (after marginalizing over all other parameters): the eXTP/SFA measurements are in red and the NICER/XTI measurements are in blue. eXTP can roughly halve the uncertainty of the NICER measurement with the same exposure time. Fig. 11 also shows the eXTP/SFA measurements of a putative 300 ks observation of the same source (in green). A longer exposure time can clearly improve the constraints on possible deviations from the Kerr geometry. Unlike spin measurements, where we want to measure the spin of as many black holes as possible, in the case of tests of GR we can properly select the most suitable sources, where systematic uncertainties are under control and subdominant with respect to the statistical ones. It is indeed plausible to image that either GR works or does not, regardless of the source.
@@ 3 Exploring Accretion Physics in Extreme Gravity
A great number of efforts have been made towards establishing the intrinsic connection between black hole X-ray binaries and AGN, such as the fundamental plane of black hole activity [147, 148] and the unified scheme for black hole accretion (e.g. [149]), in which it is proposed that different types of AGN are determined by Eddington ratio (Eddington rate scaled mass accretion rate) and the viewing angle. As for BHXRBs, the basic picture is in place but many key questions still need to be settled, such as the structure of the disk and corona, the physical mechanism of the production and evolution of AGN QPOs, the origin of quasi-periodic eruptions (QPEs) and the implications for AGN feedback. The latter question has wide relevance, because the process of accretion on to black holes plays an important role in shaping the Universe. The energy released in a small region (e.g., a few tens of gravitational radii) near the black holes in AGN can cause significant feedback which affects the evolution of their host galaxies.
3.1 The geometry and dynamics of the accretion flow in stellar mass black holes
@@ 3.1.1 Disk-corona time delays
In addition to spectral analysis, timing analysis has been employed to study the accretion geometry and its evolution. Low- and high-frequency noise in the power spectrum of the X-ray fast variations carries information about the accretion flow [154]. In some X-ray binaries, it has been observed that at low frequencies, the hard X-ray variation lags behind the soft X-ray (hard delay), and the time delay decreases with decreasing optical variation time scale. These ‘hard lags’ are thought to arise from perturbations in the accretion rate at different radii of the accretion disk propagating inward [155]. In contrast, at high variability frequencies (typically >1>1 Hz), soft X-ray variations are observed to lag behind those in hard X-rays (soft delay or ‘soft lags’). This effect may originate from the time difference corresponding to the irradiation of the accretion disk by the thermal accretion flow, i.e., the X-ray reverberation signal.
@@
[…]
where a and b are the semi-major and semi-minor axis of the hyperboloid. $\textit{h}_{\rm t}$ and $\textit{h}_{\rm b}$ represent the heights of the top and bottom boundaries, respectively. The reference coordinate system is Cartesian, with the black hole at zero, where a geometrically thin and optically thick disk is situated in the xy plane, depicted by the blue solid line in Fig. 12. The disk has a truncation radius of $10R_{g}$ and an external radius of $500R_{g}$. Two distinct geometric types (disk-like and jet-like) correspond to varying parameters of the hyperboloid (see Fig. 12). The disk-like corona is wider than it is tall, resembling a radially extended disk, while the jet-like corona appears the opposite, resembling a jet.
@@
Figure 12: Cross-section of the hyperbolic corona in the black hole frame. a and b are the semi-major and semi-minor axis of the hyperboloid. For disk-like corona (black solid line): a = 10, b = 2, upper boundary height $\textit{h}_{\rm t}=30$, lower boundary height $\textit{h}_{\rm b}=6$. For jet-like corona (red solid line): a = 120, b = 200, $\textit{h}_{\rm t}=180$, $\textit{h}_{\rm b}=2$. The units of all numbers are $R_{g}$.
@@
For the light curve, the count rates in each energy bin were multiplied by the mean effective area for the corresponding energy bin. The count rates within the specified energy bands (0.5-1.0 keV and 2.0-5.0 keV) were summed to produce the light curve. Using the Python library Stingray v2.0.0 [163, 164, 165], the time lag-frequency between the 0.5-1.0 keV and 2.0-5.0 keV bands was computed. Fig. 14 presents the results, where a positive lag indicates a hard lag, meaning the hard photons are delayed relative to the soft photons. The simulation results in Fig. 13 and Fig. 14 clearly indicate that, while the energy spectra of the disk-like and jet-like corona are indistinguishable, the frequency-dependent lag reveals significant differences.
@@ 3.1.2 Reflection reverberation
Figure 15: Spectral model and simulated eXTP-SFA lag vs. energy spectra for a 1 Crab hard state BHXRB, showing that two distinct inner geometries for disk inner radius ($r_{\rm in}$ and source height ($h$) can be easily distinguished with the lags. The upper panel shows the assumed spectral models with disk emission (dotted lines) set to have similar disk temperature, making the overall spectral energy distributions nearly identical. The lower panel shows the simulated lag data. The simulation assumes 100 ks exposures and shows the lags obtained for the 10-20 Hz frequency range, where evidence for soft reverberation lags has been seen by NICER and where reverberation delays are expected to dominate over propagation delays, which are washed out on these time-scales. The simulation uses the reltrans [83, 84] lamp-post model for reverberation lags. The lag model curves are shown with the smooth solid lines and data and model are set to zero lag at 2 keV so the shapes can be easily compared.
@@ 3.1.3 Polarization measurements of coronal geometry
These questions are not expected to be solved with the IXPE instrument, given the fast evolution of the transient black hole X-ray binaries, their typically dim fluxes in the initial outburst stages and the requirement of polarization measurement with sub-per cent accuracy. To get to a precision of 0.2% on polarization degree, required to constrain the outflow parameters, a generic 100 mCrab hard state black hole X-ray binary source would need to be observed by IXPE for more than a week (accounting for the $\sim$50% effective exposure time). This time is comparable to the characteristic timescales of the outburst evolution, preventing precise measurements of the geometry at distinct times, and a clear understanding of geometric evolution. In contrast, as shown in Fig. 16, eXTP/PFA brings the exposure time down to about a day, hence becoming uniquely suited to track the evolution of accretion geometry for black hole transients.
Figure 16: Simulated eXTP/PFA data for a generic BHXRB source at the initial stages of an outburst. The binary is assumed to have an inclination of $i=30^{\circ}$, with an incident spectrum characterized by a photon index $\Gamma=1.6$ in the 2–8 keV range and an unabsorbed flux of 100 mCrab. The model PD corresponds to a slab corona geometry with an outflow relative velocity of $\beta=0$ (static, black solid line) and $\beta=0.4$ (mildly relativistic, black dashed line), based on [101]. An eXTP exposure of 100 ksec is sufficient to distinguish between these alternatives.
@@ 3.1.4 Low-frequency QPOs as a disk-corona probe
The geometry of the corona surrounding black holes is not static during outbursts. Recent studies propose that the coronal geometry undergoes dynamic evolution, particularly during spectral state transitions from the hard to soft state [e.g. 171, 172]. High-cadence observations with the eXTP mission will be uniquely positioned to probe this evolution through its unprecedented combination of simultaneous X-ray spectral, timing, and polarimetric measurements. Furthermore, eXTP’s capacity for prolonged exposure significantly enhances the likelihood of capturing rapid changes of corona during state transitions [e.g. 173, 174, 175].
@@
3.2 Mapping the inner regions of supermassive black holes
@@ 3.2.1 X-ray reverberation mapping of the inner regions
As for accreting stellar-mass black holes, the measurement of energy-dependent time delays, i.e., lag-energy spectra, are useful to test detailed models for the innermost emitting regions of AGN, such as propagating fluctuations and reverberation [e.g., 167]. However, the spectral-timing measurements require observations with very high count rates, so can only be done for a handful of bright AGNs with current instruments. Furthermore, for past and current proportional-counter instruments with large collecting area at Fe K energies (e.g. RXTE, ASTROSAT), the spectral resolution is poor, which prevents detailed investigation of reverberation properties in the Fe K band. With its large effective area and CCD-quality spectral-resolution in the 0.5–10 keV band, eXTP/SFA will provide much better signal-to-noise measurements of the lag-energy spectra, at sufficient resolution to resolve Fe K reverberation signatures.
@@
Given the large effective area of eXTP-SFA, we can perform X-ray reverberation mapping of both the continuum and Fe K line for the brightest nearby intermediate-mass black holes (IMBHs, 102-106 M), such as NGC 4395. IMBHs serve as a crucial population bridging stellar-mass black holes and supermassive black holes (SMBHs). Measuring X-ray time lags in IMBHs is essential for filling the gap between these two mass regimes and for establishing a global scaling relation between the continuum/emission-line lags and black hole mass [167]. These measurements also provide key insights into the corona-disk-jet connection across different accretion modes [199]. Furthermore, recent IXPE observations [200] suggest that the corona geometry extends along the same direction as the accretion disk, challenging the simple lamp-post model assumption. eXTP offers a unique opportunity to test this corona illumination and disk reprocessing scenario, providing new constraints on the physics of accretion and black hole environments.
@@ 3.2.2 AGN non-thermal superflare from an orbiting hotspot at the SMBH inner radii
Recently, Very Large Telescope Interferometer (VLTI) observed several near-infrared superflares which might be from hot spots moving around Sgr A* at only 7–10 gravitational radii from the central black hole [204]. It is believed that these hot spots are likely due to relativistic magnetic reconnection events and they might be also the sources of X-ray flares which exhibit a double-peak structure in the in the light curve of the total X-ay flux due to their orbital motion. Here we simulate the X-ray light curve of these hot spots observed by eXTP-SFA.
@@
Fig. 18 (left) shows a simulation of the light curve produced by a hot spot orbiting a black hole for one period, as observed by eXTP-SFA in the 2–10 keV energy band. We simulate a 4.2 ks observation with an average flux of $10^{-11}$ erg cm-2 s-1. The black hole mass is assumed to $4.14\times 10^{6}\,M_{\odot}$ and black hole spin is 0.998. The orbital radius of the hot spot is $r_{\mathrm{spot}}=10\,R_{g}$. Its radiation is assumed as a power law spectrum with index $\Gamma=2$. The observing inclination is assumed to be $75^{\circ}$. There is an obvious double-peak structure in the light curve, with a time interval of 500 s between the two peaks. In the lower subpanel, we show the ratio between the data and the best fit model. We obtain a good fit and recover the input parameters.
Fig. 18 (right) shows the error contours (1, 2, and 3 $\sigma$ in red, green and blue, respectively) of hot spot orbital radius and observing inclination. The orbital radius of the hot spot can be measured to a precision $3\,R_{g}$ at $3\sigma$ confidence, while the inclination is measured to a precision of $1^{\circ}$. The technique also provides us with a way to measure the black hole spin, if the mass of the black hole is measured independently (e.g. via optical reverberation mapping [205]). This would apply to any hot spot radius and hence is independent of whether or not the disk is assumed to extend to the ISCO, as is the case for spin constraints from spectral fitting. Thanks to its unprecedented throughput, eXTP will be able to perform such measurements for any AGN with X-ray flux above 1 mCrab.
@@ 3.2.3 Quasi Periodic Oscillations from SMBHs
AGN, powered by accretion on to SMBHs at the center of of galaxies, are thought to be scaled-up versions of BHXRBs [206, 207, 208, 209]. A compelling line of evidence for this postulation is the striking similarity in the variability of the X-ray radiation between AGN and BHXRBs [e.g. 210, 211, 212, 213, 214]. The power spectral densities (PSDs) of both can be well described as red-noise. However, the characteristic QPOs, which have been observed in the X-ray light curves of dozens of BHXRBs [e.g. 215, 46], are rarely detected in AGN. More detections of QPOs in SMBHs will strengthen the connection between AGN and BHXRBs. Also, the properties of QPOs in AGN can provide a powerful probe to diagnose the inner structure of the X-ray emitting region around SMBHs.
@@
So far, only a few possible QPOs have been found in accreting systems with SMBHs: the Narrow-Line Seyfert 1 (NLS1) galaxies RE J1034+396 [222], 1H 0707-495 [223], ESO 113-G010 [224], Mrk 766 [225], MCG-06-30-15 [226], MS 2254.9–3712 [227], the ultrasoft AGN candidate 2XMM J123103.2+110648, as well as the tidal disruption events (TDE) Swift J164449.3+573451 [228], etc. But the reliability of some cases are still under debate [e.g. 229]. Considering the the dependence of the timescale on the black hole mass [46, 27], it is not hard to understand the lack of QPO detections in SMBHs. More QPOs in SMBHs will be found if more high-quality observations with long, continuous, uninterrupted exposure can be taken.
[…]
Figure 19: Left panel: the PSD of RE J1034+396 observed with XMM-Newton. The blue line represents the best-fit power law, the orange line represents the PSD model with an added Lorentzian profile, and the red dashed line represents the 99.99% confidence level. Middle and right panels: the examples of simulated PSDs observed with SFA, with flux of $1\times 10^{-11}\,\rm erg\,s^{-1}\,cm^{-2}$ and $2.5\times 10^{-13}\,\rm erg\,s^{-1}\,cm^{-2}$, respectively.